OpenStudy (anonymous):

Write the equation for the following relation. D = {(x, y): (3, 9), (5, 25), (-2, 4), . . .}

hartnn (hartnn):

did u try to find the relation ? how is 3 related to 9 ? 5 related to 25 ?

OpenStudy (anonymous):

as a hint: look at all y/x... need more explanation?

OpenStudy (anonymous):

y1/x1 = 9/3 = 3 y2/x2 = 25/5 = 5 y3/x3 = 4/-2 = -2 . . . right?

hartnn (hartnn):

no, its x^2=y

OpenStudy (anonymous):

yes... of course it is that... just gave hint to reach it,... not giving complete answer.